OpenStudy (moongazer):

What are the points of intersection of y = 8 and f(x) = x(x+2)? Please answer clearly and explain the solution clearly and easy to understand. Thanks! (guaranteed medal for a good answer)

OpenStudy (anonymous):

8=x(x+2) 8=x^2+2x x^2+2x-8=0 (x+4)(x-2)=0 x=-4 x=2